OpenStudy (anonymous):

Can someone PLEASE check my answer? Given f(x) = (2x - 2)/4, solve for f^−1(3). I got 7.

OpenStudy (mathmate):

Yes, first find \(f^{-1}(x)\) y=(2x-2)/4 x=(2y-2)/4 (swap x, y) 4x=2y-2 (solve for y) y=2x+1 \(f^{-1}(x)=2x+1\) \(f^{-1}(3)=2(3)+1=7\) Yes your answer is correct!

OpenStudy (anonymous):

thank you :)

OpenStudy (mathmate):

you're welcome! :)